the sum of the present ages of sheila and kristine is 42 years. in 9 years, sheila will be 4 years older than 2 times as old as kristine was 2 years ago. find their present ages.
Answer by lwsshak3(11628)

the sum of the present ages of sheila and kristine is 42 years. in 9 years, sheila will be 4 years older than 2 times as old as kristine was 2 years ago. find their present ages.
**
let x=sheila's present age
42-x=kristine's present age
..
In 9 years:
x+9=sheila's age
..
2 years ago:
42-x-2=40-x=kristine's age
..
x+9=2(40-x)+4
x+9=80-2x+4
3x=75
x=25
42-x=17
sheila's present age=25
kristine's present age=17
..